# Fine-Grained Assertion-Based Test Selection

### v1

@@ IV-C Procedure and Results
LogicNG 179.06 (32.78) 175.77 3.29 12.08 (2.20) 9.05 3.03 744.54 767.02 641.25 632.86 18.60 18.86 -0.26
Ave. 74.32 75.32 -0.99 6.83 6.28 0.55 146.18 152.20 412.16 418.67 15.11 15.65 -0.54
To evaluate the performance of Selertion, we measure the offline end-to-end time (AE time), analysis time (A time), collection time (C time) for dependency collection which runs in a separate process, as well as the initial overhead, respectively. In addition, we disable the finest-grained selection component of Selertion to assess the impact of selecting test entities at the assertion level. After disabling the component, the granularity of selection becomes test method level. To have a safe analysis, selection at the test method level still requires instrumenting the test code at statement level to inspect whether there exists any test method that depends on other tests. As before, the test classes that contain test methods calling other tests are selected at test class level along with those using inheritance or parameterized tests. For the AE time, we already presented the results for Selertion in Table III. To compare with the method level selection, we present the results in seconds and percentages in Table IV under End-to-end Time (AE). We report the difference for AE time between Selertion and the method-level variant. Our analysis reveals that, when comparing assertion-level selection to method-level selection, there is some variation in end-to-end time across different subjects. For subjects like commons-exec, OpenTripPlanner, accumulo-core, and commons-pool, assertion-level selection consistently requires less end-to-end time. However, for the remaining subjects, assertion-level selection tends to demand slightly more time, with differences ranging from 0.1 second (as seen in commons-net) to 3.29 seconds (observed in LogicNG) on average. Notably, the overall average for time difference is -0.99 seconds, indicating that assertion-level selection is, on average, faster. To gain further insights from our results, we group the 11 subjects into two categories: the first one consisting of subjects with lower average times than method-level selection, and the second one including subjects with higher times. We conduct a Paired T-Test ($\alpha=0.1$) for these two groups, aiming to compare the results between Selertion and its coarser-grained variant. Selecting $\alpha=0.1$ as the significance level enhances the sensitivity of our test [42]. The outcomes of the analysis indicate that, for the first group, Selertion is significantly faster than the method-level selection. In contrast, for the second group, which comprises subjects with higher end-to-end times, there is no statistical difference in offline end-to-end running time between Selertion and its coarser-grained variant. Therefore, we can confidently conclude that Selertion is either significantly faster or equivalent to its method-level variant. The analysis time (A) is positively correlated to the size of code changes, because large code changes require more time for change computation. For example, because the revisions of tika-parsers have relatively large code changes, as shown in Figure 4 where the boxes for the subject are higher than the others, tika-parsers has the longest analysis times of 18.93 seconds. However, for the long-running subjects, the analysis time only takes a small portion of the overall time. Tika-parsers has 18.93 seconds of analysis time on average which is only 10.48% of the RetestAll. In contrast, asterisk-java has 3.84 seconds analysis time which is much shorter than 18.93 seconds, but the analysis time is 11.72% of the RetestAll. That might also explain why Selertion works better on long-running subjects. To compare with its method-level variant, from the results in the columns under Analysis Time (A) in Table IV, the analysis time for Selertion is more than method-level selection for each subject. Overall, it takes 0.55 second more on average than method-level selection for the analysis phase. Besides the mean values reported in Table IV, we also calculate the Paired T-Test. The p-value is much less than $\alpha$ (i.e., $p<\alpha$, where $\alpha=0.1$); thus, we conclude that fine-grained assertion selection requires more analysis time than method-level selection, as expected. For our implementation, the collection phase is in a separate process, meaning developers can obtain the RTS results first and there is no need to wait for the data collection to finish. But it is still necessary to measure collection time to evaluate the performance. The results of collection time are shown under Collection Time (C) in Table IV. Considering all the 11 subjects as a whole, on average, the collection time is 146.18 seconds for Selertion, while 152.20 seconds for method-level selection. However, based on the Paired T-Test, the collection time for method level selection has no statistical difference with Selertion. The initial overhead includes code instrumentation, backward assertion slicing and dependency collection at the initial run. In table IV, we present the results under Initial Overhead. For assertion-based selection, the time consumption is from 34.26 seconds for asterisk-java to 887.99 seconds (14.8 mins) for tika-parsers. Similarly, the result of Paired T-Test indicates there is no statistical difference for method level selection in terms of initial overhead. Because the initial overhead is just a one-time requirement, it is acceptable compared to the execution time saved by our tool over time. Besides the time components, we also compare the selected assertion ratio between Selertion and the method-level variant. The results are shown in the fifth macro-column in Table IV. Overall, fine-grained test selection will select less assertions than method-level selection. On average, Selertion chooses 0.54% fewer assertions per revision for each subject, which translates to 14 fewer assertions in absolute numbers. Similarly, The p-value of Paired T-Test for selected assertion ratio indicates that the fine-grained selection is significantly better than method-level selection in terms of selected assertion ratio. After comparing with the variant by disabling fine-grained assertion selection, although Selertion requires more analysis time, its end-to-end time, collection time and initial overhead are either less or equivalent to the coarse-grained variant. This indicates that fine-grained selection generally does not consume more time or result in a performance loss. Moreover, fine-grained assertion-level selection is either significantly faster or equivalent to its method-level variant and exhibits the ability to select fewer assertions with higher precision across all subjects. Therefore, we can confidently conclude that fine-grained assertion-level selection surpasses method-level selection in terms of overall performance and precision.
@@ V Discussion
Fine-Grained Selection Safety. In terms of production code analysis, we follow the design decisions of prior works HyRTS [12] and Ekstazi [9]; both argue they are safe in their test selection. Our analysis granularity for production code is a combination of method and class level, because we transform changes outside method bodies to coarse-grained class-level changes. It has been proven by Zhang [12] that this kind of change transformation cannot introduce new safety issues. Thus, if there are any potential safety issues in our technique, they may derive from the fine-grained instrumentation and selection for test code. Our assertion slicing component (see Section III-B) determines whether a test can be sliced into assertion slices. We have performed many experiments to make sure the assertion slices are executable without breaking the original test methods. To guarantee a safe selection, we perform specific analysis (see Section III-C) to discreetly identify tests that are unfeasible for finer-grained selection including both test method and assertion level and take a conservative selection approach (see Section III-D). For tests containing statements out of our analysis scope we also take a conservative approach to be safe. For instance, test statements that call third-party libraries such as Thread.sleep, are not safe to be selected at assertion level, because the statement Thread.sleep will not be included in any assertion slices. We downgrade such tests to method-level selection to be safe.
Test Suite Effectiveness We employed Major [36] to randomly generate mutants across different methods in the subjects and compared the effectiveness of the reduced test suite by Selertion with the original test suite in eliminating these mutants. In all subjects, the mutants killed by Selertion’s reduced test suite align with those killed by the original test suite. For example, for commons-math with the most test cases in our evaluation, Selertion achieves an identical mutation score of 47% when compared to the original test suite. These empirical findings affirm that our fine-grained selection approach maintains the same test effectiveness as observed in the original test suite, and thus is safe.
@@
Source code vs. Bytecode. Different from the other tools that identify changes using Java bytecode, we focus on changes at source code level. There are some scenarios that may lead to completely different results. For example, if the program revision is to upgrade the Java version without any real source code changes, our tool will not select any tests, however, because the bytecode is changed completely, the other tools will select the whole test suite. Moreover, some modifications of source code do not impact the bytecode, such as changing a+=1 to a=a+1; here Selertion selects tests related to the source code changes, however, the other tools will fail to select any tests.
@@ VI Related Work
Regression test selection (RTS) has been a widely studied topic, with various techniques ranging from basic-block-level to module-level granularity reviewed in several surveys [43, 44, 45, 7, 8]. The basic-block-level RTS, introduced by Rothermel and Harrold for C programs and later extended to Java [46, 39], evolved into coarser-grained methods to improve efficiency. Notable among these are FaultTracer and Chianti at method level [11, 10], and Ekstazi at file level [9], with Ekstazi demonstrating superior performance in testing time and scalability. Further research explored coarser granularities, like file-level and module-level RTS [47, 48]. The trade-off between precision and efficiency led to hybrid approaches, such as Orso et al.’s two-phase RTS [41] and Zhang’s HyRTS, which combines method and class level analysis [12]. Alongside dynamic RTS, static RTS techniques have been explored, though they face challenges with safety and practicality [49, 40, 50]. Current RTS research primarily focuses on the granularity of production code analysis, typically selecting test entities at class or method level. Building on the concept of checked coverage [51] and its correlation with test suite effectiveness [14], our work innovates by selecting tests at the assertion level, treating each assertion and its slice (i.e., associated statements) as the smallest selection unit. Research on breaking down test cases for assertion-related studies includes Chen et al.’s assertion-aware test reduction [52], Vahabzadeh et al.’s fine-grained test case minimization [34], Fang et al.’s assertion fingerprints for identifying similar test cases [53], and Xuan et al.’s work on enhancing dynamic analysis and test purification [54, 55]. However, none of these specifically target RTS at the assertion level, making our approach a novel contribution to the field.

### v2

@@ IV-C Procedure and Results
LogicNG 179.06 (32.78) 175.77 3.29 12.08 (2.20) 9.05 3.03 744.54 767.02 6216.11 6207.72 18.60 18.86 -0.26
Ave. 74.32 75.32 -0.99 6.83 6.28 0.55 146.18 152.20 918.96 925.47 15.11 15.65 -0.54
To evaluate the performance of Selertion, we measure the offline end-to-end time (AE time), analysis time (A time), collection time (C time) for dependency collection which runs in a separate process, as well as the initial overhead, respectively. In addition, we disable the finest-grained selection component of Selertion to assess the impact of selecting test entities at the assertion level. After disabling the component, the granularity of selection becomes test method level. To have a safe analysis, selection at the test method level still requires instrumenting the test code at statement level to inspect whether there exists any test method that depends on other tests. As before, the test classes that contain test methods calling other tests are selected at test class level along with those using inheritance or parameterized tests.
For the AE time, we already presented the results for Selertion in Table III. To compare with the method level selection, we present the results in seconds and percentages in Table V under End-to-end Time (AE). We report the difference for AE time between Selertion and the method-level variant. Our analysis reveals that, when comparing assertion-level selection to method-level selection, there is some variation in end-to-end time across different subjects. For subjects like commons-exec, OpenTripPlanner, accumulo-core, and commons-pool, assertion-level selection consistently requires less end-to-end time. However, for the remaining subjects, assertion-level selection tends to demand slightly more time, with differences ranging from 0.1 second (as seen in commons-net) to 3.29 seconds (observed in LogicNG) on average. Notably, the overall average for time difference is -0.99 seconds, indicating that assertion-level selection is, on average, faster. To gain further insights from our results, we conduct a Paired T-Test ($\alpha=0.05$) to compare the performance of Selertion with its coarser-grained variant. The analysis reveals that, for the first group of subjects, including tabula-java, accumulo-core, and commons-pool, Selertion is significantly faster than the method-level selection, with a p-value of 0.041, which is lower than $\alpha$. In contrast, for the second group, which includes the remaining eight subjects, there is no statistically significant difference in the offline end-to-end running time between Selertion and its coarser-grained variant. Therefore, we can confidently conclude that Selertion is either significantly faster or equivalent to its method-level variant.
The analysis time (A) is positively correlated to the size of code changes, because large code changes require more time for change computation. For example, because the revisions of tika-parsers have relatively large code changes, as shown in Figure 4 where the boxes for the subject are higher than the others, tika-parsers has the longest analysis times of 18.93 seconds. However, for the long-running subjects, the analysis time only takes a small portion of the overall time. Tika-parsers has 18.93 seconds of analysis time on average which is only 10.48% of the RetestAll. In contrast, asterisk-java has 3.84 seconds analysis time which is much shorter than 18.93 seconds, but the analysis time is 11.72% of the RetestAll. That might also explain why Selertion works better on long-running subjects. To compare with its method-level variant, from the results in the columns under Analysis Time (A) in Table V, the analysis time for Selertion is more than method-level selection for each subject. Overall, it takes 0.55 second more on average than method-level selection for the analysis phase. Besides the mean values reported in Table V, we also calculate the Paired T-Test. The p-value is 0.015 which is much less than $\alpha$ (i.e., $p<\alpha$, where $\alpha=0.05$); thus, we conclude that fine-grained assertion selection requires more analysis time than method-level selection, as expected.
In our implementation, the collection phase operates as a separate process, allowing developers to obtain RTS results without needing to wait for the data collection to complete. However, it is still important to measure collection time to evaluate performance. The results for collection time are presented under Collection Time (C) in Table V. The collection time varies among subjects. For some subjects, including asterisk-java, commons-net, accumulo-core, commons-exec, tabula-java, and commons-pool, the collection time is either shorter than or comparable to the AE time. However, for a subject like LogicNG, the collection time is significantly longer. Upon manual inspection of the LogicNG codebase, it was found that the test cases include heavy use of for loops. We plan to improve the instrumentation for test cases with extensive loop executions in the future. Nonetheless, our implementation decouples the collection phase from the test selection approach, allowing developers to control when and whether to trigger the collection process. This process runs in the background on a separate copy of the codebase, ensuring it does not interfere with the developer’s work on the original codebase. Considering all the 11 subjects as a whole, on average, the collection time is 146.18 seconds for Selertion, while 152.20 seconds for method-level selection. However, based on the Paired T-Test, the collection time for method level selection has no statistical difference with Selertion.
The initial overhead includes code instrumentation, backward assertion slicing and dependency collection at the initial run. In Table V, we present the results under Initial Overhead. For assertion-based selection, the average time consumption is 918.96 seconds (15.3 min). Similarly, the result of Paired T-Test indicates there is no statistical difference for method level selection in terms of initial overhead. Because the initial overhead is just a one-time requirement, it is acceptable compared to the execution time saved by our tool over time.
Besides the time components, we also compare the selected assertion ratio between Selertion and the method-level variant. The results are shown in the fifth macro-column in Table V. Overall, fine-grained test selection will select fewer assertions than method-level selection. On average, Selertion chooses 0.54% fewer assertions per revision for each subject, which translates to 14 fewer assertions in absolute numbers. The p-value of Paired T-Test is 0.006 for selected assertion ratio indicates that the fine-grained selection is significantly better than method-level selection in terms of selected assertion ratio.
After comparing with the variant by disabling fine-grained assertion selection, although Selertion requires more analysis time, its end-to-end time, collection time and initial overhead are either less or equivalent to the coarse-grained variant. This indicates that fine-grained selection generally does not consume more time or result in a performance loss. Moreover, fine-grained assertion-level selection is either significantly faster or equivalent to its method-level variant and exhibits the ability to select fewer assertions with higher precision across all subjects. Therefore, we can conclude that fine-grained assertion-level selection surpasses method-level selection in terms of overall performance and precision.
@@ V Discussion
Fine-Grained Selection Safety. In addition to the analysis in Section IV-C4, where we leveraged mutation testing to evaluate the test effectiveness across all subjects, we manually introduced real bugs from the Defects4J v2.0.1 [45] database. Upon inspection, we found that the test suites for certain projects in Defects4J, such as Chart and JacksonCore, are still using JUnit3, while our tool is designed to support JUnit4. As a result, we chose to assess our tool with real bugs in commons-math, the project common to both our subject systems and Defects4J. Selertion successfully selected the same triggered tests as those documented in Defects4J for all the bugs. Regarding our approach to production code analysis, we follow the design principles of prior works, such as HyRTS [12] and Ekstazi [9]; both argue they are safe in their test selection. Our analysis granularity for production code is a combination of method and class level, because we transform changes outside method bodies to coarse-grained class-level changes. It has been proven by Zhang [12] that this kind of change transformation cannot introduce new safety issues. Additionally, for inheritance in production code, since the inheritance relationship is captured during code instrumentation, if changes in parent class A impact child class B, all test entities related to both classes A and B will be selected.
@@
Thus, if there are any potential safety issues in our technique, they may derive from the fine-grained instrumentation and selection for test code. Our assertion slicing component (see Section III-B) determines whether a test can be sliced into assertion slices. We have performed many experiments to make sure the assertion slices are executable without breaking the original test methods. To guarantee a safe selection, we perform specific analysis (see Section III-C) to discreetly identify tests that are unfeasible for finer-grained selection including both test method and assertion level and take a conservative selection approach (see Section III-D). For tests containing statements out of our analysis scope we also take a conservative approach to be safe. For instance, test statements that call third-party libraries such as Thread.sleep, are not safe to be selected at assertion level, because the statement Thread.sleep will not be included in any assertion slices. We downgrade such tests to method-level selection to be safe.
@@
Source code vs. Bytecode. Different from the other tools that identify changes using Java bytecode, we focus on changes at source code level. We would like to mention that there are some scenarios that may lead to completely different results. For example, if the program revision is to upgrade the Java version without any real source code changes, our tool will not select any tests, however, because the bytecode is changed completely, the other tools will select the whole test suite. In cases where configuration changes occur without any real source code changes, developers might choose to execute the entire test suite without relying on an RTS tool. Moreover, some modifications of source code do not impact the bytecode, such as changing a+=1 to a=a+1; here Selertion selects tests related to the source code changes, while the other Bytecode analysis based tools will not select any tests. We believe that our fine-grained, source code analysis-based tool is more suitable for developers to verify their source code changes in daily tasks, whereas bytecode-based tools are better suited for use in a continuous integration environment.
@@ VI Related Work
Regression test selection (RTS) has been a widely studied topic, with various techniques ranging from basic-block-level to module-level granularity reviewed in several surveys [46, 47, 48, 7, 8]. The basic-block-level RTS, introduced by Rothermel and Harrold for C programs and later extended to Java [49, 40], evolved into coarser-grained methods to improve efficiency. Notable among these are FaultTracer and Chianti at method level [11, 10], and Ekstazi at file level [9], with Ekstazi demonstrating superior performance in testing time and scalability. Further research explored coarser granularities, like file-level and module-level RTS [50, 51]. The trade-off between precision and efficiency led to hybrid approaches, such as Orso et al.’s two-phase RTS [42] and Zhang’s HyRTS, which combines method and class level analysis [12]. Alongside dynamic RTS, static RTS techniques have been explored, though they face challenges with safety and practicality [52, 41, 53].
Current RTS research primarily focuses on the granularity of production code analysis, typically selecting test entities at class or method level. Building on the concept of checked coverage [54] and its correlation with test suite effectiveness [14], our work innovates by selecting tests at the assertion level, treating each assertion and its slice (i.e., associated statements) as the smallest selection unit.
Research on breaking down test cases for assertion-related studies includes Chen et al.’s assertion-aware test reduction [55], Vahabzadeh et al.’s fine-grained test case minimization [34], Fang et al.’s assertion fingerprints for identifying similar test cases [56], and Xuan et al.’s work on enhancing dynamic analysis and test purification [57, 58]. However, none of these specifically target RTS at the assertion level, making our approach a novel contribution to the field.